OpenStudy (anonymous):

Which of these are true or false? 1. sec^2(x) - 1 = tan^2(x) 2. sin2(x) / 2 = sin(x) 3. sin(x)+cos(x) / sin(x) = cos(x) 4. sin^2(x) = cos^2(x) -1 5. cot(x) + 1 = csc(x) 6. 2sin(x) = sin(x) + sin(x)

OpenStudy (anonymous):

Fundamentally\[ \sin^2(x)+\cos^2(x)=1 \]Divide both sides by \(\cos^2(x)\) and you get\[ \tan^2(x) + 1 = \sec^2(x) \]Divide both sides by \(\sin^2(x)\) and you get: \[ 1+\cot^2(x) = \csc^2(x) \]

OpenStudy (anonymous):

Then you can just check if the other formula seem to match up with these formula